Find the expected value

E(X),

the variance

Var(X)

and the standard deviation

σ(X)

for the density function. (Round your answers to four decimal places.)

f(x) =

1
x

on

[1, e]

E(X) =
Var(X) =
σ(X) =
Answer #1

Var(X) = E(X^2) - E(X)^2

= 3.19452 - 1.71828^2

= 0.2420

sd= sqrt(Var(X) = sqrt( 0.2420 ) = 0.491969

Please give me a thumbs-up if this helps you out. Thank you! :)